OpenStudy (anonymous):

log3(xy)^3-log3xy= 3log3????

OpenStudy (across):

I'm not sure I understand the equation. Is it\[\log_{3}{(xy)^3}-\log_{3}{(xy)}=3\log_{10}{(3)}?\]

OpenStudy (anonymous):

yes

OpenStudy (across):

First of all, do you think we can simplify the left-hand-side? Perhaps using this property\[\log(a)-\log(b)=\log(\frac{a}{b})?\]

OpenStudy (anonymous):

no, the 3log3 was my answer to the question

OpenStudy (across):

Well, the LHS of the expression is a function of x and y whilst the RHS is a constant. How can that be?

OpenStudy (anonymous):

what is LHS?

OpenStudy (across):

left-hand-side

OpenStudy (anonymous):

left hand side

OpenStudy (anonymous):

I have never seen these problems in my life, I only started Logarithms today.

OpenStudy (across):

\[\log_{3}(xy)^3-\log_{3}(xy)=\log_{3}(x^3y^3)-\log_{3}(xy)=\log_{3}(\frac{x^3y^3}{xy})=?\]

OpenStudy (anonymous):

log x2 y2?

OpenStudy (anonymous):

base 3

OpenStudy (across):

That's correct.

OpenStudy (anonymous):

and that cam be simplified to 2log3(xy) right??

OpenStudy (across):

Yep:\[\log_{3}{(x^2y^2)}=\log_{3}{(xy)^2}=2\log_{3}{(xy)}.\]

OpenStudy (across):

It's all about the properties...

OpenStudy (anonymous):

thank you ms MIT